OpenStudy (anonymous):

A pharmacist is to prepare 15ml of special eye drops for a glaucoma patient. The eye drop solution have a 2% active ingredient, but the pharmacist has only 10% solution and 1 % solution in stock. How much of each type of solution must be used to fill the prescription?

OpenStudy (anonymous):

The two solutions together should be 15 ml.So let there be x ml of 10% solution and y ml of 1% solution. Now x+y=15 Also concentrations should match, so, 0.1x + 0.01y =.02*15 Solving the two simultaneous equations we get, 0.09x +0.15 =0.3 x=1.666ml of 10% solution and the rest is 1% solution